OpenStudy (anonymous):

Solve the inequality algebraically. 5 < 6 X+3

OpenStudy (nikato):

Can u rewrite that more clearer?

OpenStudy (anonymous):

sure hold on!

OpenStudy (anonymous):



OpenStudy (nikato):

Ok, multiply both sides by x+3

OpenStudy (anonymous):

ok can u show me i am sorry i am not good at this stuff! sorry

OpenStudy (nikato):

(X+3) 5. < 6 (x+3) -------------- (x+3)

OpenStudy (nikato):

The x+3 on the left cancel out. So ur left with 5<6(x+3)

OpenStudy (nikato):

Now distribute the right side

OpenStudy (nikato):

Now Subtract 18 on both sides And divide both sides by 6

OpenStudy (anonymous):

\[\frac{ x+3 }{5 }>\frac{ 1 }{6},x+3>\frac{ 5 }{6 },x>\frac{ 5 }{ 6 }-3,x>\frac{ -13 }{ 6 }\]

OpenStudy (anonymous):

thanks yall!(:

OpenStudy (anonymous):

when you multiply by x+3 ,you can say nothing about inequality. x+3 may be positive or negative,so avoid multiplying by x+3